In order to measure the volume of a piece of apparatus, Max filled a 750mL flask with 700mm Hg pressure of gas, and then expanded it into the apparatus. The final pressure was 224mm Hg. Calculate the total volume occupied by the gas and hence the volume of the apparatus being measured.
1
Expert's answer
2020-09-15T13:52:15-0400

Solution:

"P_{1}V_{1}=P_{2}V_{2}"

"700mmHg(750mL)=224mmHg(V_{2})"


"V_{2}= \dfrac{700mmHg(750mL)}{224mmHg}"


"V_{2}=2344mL=2.34L"


Since the original flask contained 0.75 L, the volume of the apparatus = 2.34 - 0.75 = 1.59 L